OpenStudy (anonymous):

Simplify: √p^6+√9p^6+3√p^6-√p^2

OpenStudy (anonymous):

\[\sqrt{p^6} + \sqrt{9p^6} +3\sqrt{p^6} -\sqrt{ p^2} = p^3 + 3p^3 + 3p^3 - p = 7p^3 - p\]

OpenStudy (anonymous):

What else you can do in this: Take p common: \[\huge 7p(p^2 - 1)\] Use: \[(a^2 - b^2) = (a-b)(a+b)\] \[\huge \color{green}{= 7p(p+1)(p-1)}\]

OpenStudy (anonymous):

\(\huge \color{blue}{\textbf{Welcome To Openstudy..}}\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

so my answers for the last problem for my study guide are 79^3 or 7p^3-p

OpenStudy (anonymous):

It is later one;; \[7p^3 - p..\]

OpenStudy (anonymous):

okay thank you soo much I have a few more questions if you don't mind

OpenStudy (anonymous):

√2(x-√5)

OpenStudy (anonymous):

Do it yourself.. I am giving you the formula: \[\huge \sqrt{a}(b - \sqrt{c}) = b \sqrt{a} - \sqrt{a \times c}\] Here, a = 2, b= x and c = 5 Plug in the values and solve..

OpenStudy (anonymous):

ok thanks and do I use the formula for this problem as well? 5√3 (2√3 +6√19)

OpenStudy (anonymous):

Yes the same formula.. You have to just distribute 5root(3) to the parenthesis..

OpenStudy (anonymous):

okay hopefully I get it

OpenStudy (anonymous):

would the answer for the first question be 2x sqrt2?

OpenStudy (anonymous):

sorry 2x-2 sqrt 2